OpenStudy (anonymous):

4x + 3y = 2 3x + 2y = 1 use elimination

OpenStudy (anonymous):

4x + 3y = 2 x=3n+2 y=-2(2n+1)

OpenStudy (anonymous):

show your work please so i can know how to do it please?

OpenStudy (anonymous):

3x + 2y = 1 x=2n+1 y=-3n-1